OpenStudy (ryan0716):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours: Age Ticket Price Under 3 years free 3 to 12 years $34 Over 12 years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (ryan0716):

(64 − x)34 + (103 − 64) ⋅ 45 = C (103 + x)34 + (103 −(64 + x)) ⋅ 45 = C (103 − x)34 + (103 − 64) ⋅ 45 = C (64 + x)34 + (103 −(64 + x)) ⋅ 45 = C

OpenStudy (mathmale):

What work have you done so far to identify the correct answer choice? Please share it. Note that you're working with rates (price per ticket sold) and counts (how many people in each age category there are). Rate * count produces the income from tickets sold in each case.

OpenStudy (ryan0716):

i was trying to figure out the the equation and I think that it is D

OpenStudy (anonymous):

#1 works